Calculate the pH of the resulting solution obtained by mixing 20 mL of
0.2M HCl solution and 50 mL of 0.1M NaOH solution by assuming complete
dissociation of HCl and NaOH in the solution?
 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
As both are the strong acids and bases respectively, the Resultant moles after totally gets dissociated would be => 
50*0.1 – 20*0.2 => 5 – 4 = 1 mmoles of NaOH would get remained.
Molarity would be => 1/70
Therefore, pOH = -log(1/70) = 1.84
pH = 14-1.84 => 12.16